# Assemble Your Crew: Automatic Multi-agent Communication Topology Design via Autoregressive Graph Generation

### v2

@@ Conclusion
In this work, we addressed two limitations in the design of multi-agent systems: the redundant composition and limited extensibility inherent in template graph modification-based approaches. We formulate the collaboration topology design problem as determining both the system composition (i.e., the set of agents and their roles) and its communication structure through autoregressive graph generation. To operationalize this paradigm, we introduced ARG-Designer, a novel autoregressive model that constructs a collaboration graph from scratch, conditioned on a natural language task query. Our approach empowers the model to dynamically determine the necessary number of agents, select appropriate roles from an extensible pool, and establish optimal communication links between them, resulting in a bespoke MAS topology precisely tailored to the specific demands of each task. Extensive experiments across six diverse benchmarks demonstrate that ARG-Designer consistently outperforms existing methods, achieving new state-of-the-art performance while maintaining superior token efficiency.

### v4

@@ Conclusion
In this work, we addressed redundant composition and limited extensibility in template graph modification approaches by reformulating collaboration topology design as autoregressive graph generation. We introduced ARG-Designer, a novel autoregressive model that constructs collaboration graphs from scratch, conditioned on natural language task queries. Our approach dynamically determines agent numbers, selects roles from an extensible pool, and establishes optimal communication links, creating bespoke MAS topologies tailored to specific task demands. Extensive experiments across six benchmarks demonstrate that ARG-Designer consistently outperforms existing methods, achieving state-of-the-art performance while maintaining superior token efficiency.
Acknowledgments
The work of S. Pan was partially supported by the Australian Research Council (ARC) under Grant Nos. DP240101547 and FT210100097.